# PTMPicker: Facilitating Efficient Pretrained Model Selection for Application Developers

### v1

@@ 1. Introduction
PreTrained Models (PTMs) refer to neural network models that have been previously trained on large-scale datasets and saved for future reuse. In recent years, the existing literature on PTMs discusses the core issues and risks of the PTM supply chain and hubs (Mitchell et al., 2019; Jiang et al., 2022; Castaño et al., 2023; Kathikar et al., 2023) and the detailed utilization of specific PTMs for special tasks (Schmidhuber, 2015; Antoniou et al., 2017; Dube et al., 2019; Xia et al., 2022; Qiu et al., 2020), such as molecular property prediction in chemistry. In addition to the literature on PTM supply chain risks and domain-specific applications, there is also a stream of research (Bolya et al., 2021; You et al., 2021; Zhang et al., 2023; Liu et al., 2023; Davis et al., 2023) dedicated to the identification and selection of suitable models. Specifically, Zhang et al. (Zhang et al., 2023) proposed Model Spider, which selects suitable pretrained models for specific tasks by computing and ranking the similarity between the candidate model token representation and the downstream task token representation. Model Spider better serves researchers and developers who are interested in the nuanced differences among various PTMs and seek to explore their performance across specific evaluation metrics. Liu et al. (Liu et al., 2023) introduces MLTaskKG to recommend ML/DL libraries/models for application developers based on the specifications of AI tasks specified by them. MLTaskKG primarily focuses on AI task descriptions while overlooking the limitations of recommended models or libraries, such as potential biases, safety concerns, and license compatibility.
The surge in PTMs has led to the emergence of Hugging Face (Hugging Face Company, nd; Wikipedia, ndb) as the de facto standard model hosting platform. Hugging Face, an online Git-based platform, hosted more than 543,000 diverse pretrained models at the time of our study, and the number of models continues to grow rapidly. This centralized repository could obviate the need for researchers to manually retrieve models from other heterogeneous hosting platforms. Although the platform provides model specifications, developers [^1] For simplicity, unless otherwise specified, the term developers hereafter refers to downstream application developers in this study. still face challenges in finding models that align with their requirements. In practice, application developers are often concerned not only with the intended functionality of potential target PTMs but also with aspects such as license compliance, model size, and interface usability. When searching using functional keywords or high-level task categories, an overwhelming number of candidate models are typically returned. Developers then need to devote considerable attention to manually and carefully examining and iterating all the candidate models, which is both labor-intensive and time-consuming.
@@ 2. Background & Motivation
Compared to training models from scratch, PTMs significantly reduce training costs and, when combined with model quantization (Egashira et al., 2024), enable deployment on resource-constrained devices. They also lower barriers to reuse for application developers with limited training expertise, allowing development goals to be met through inference with minimal integration effort.
@@
The emergence and widespread adoption of PTMs has attracted significant research attention, with some researchers focusing on the development and analysis of model documentation, commonly referred to as model cards. Mitchell et al. (Mitchell et al., 2019) first proposed model cards for two types of Deep Learning (DL) models: one trained to identify smiling faces in images and the other trained to detect offensive comments in natural language texts. These model cards were intended to promote model transparency among not only users but also developers and stakeholders. Going one step further, Crisan et al. (Crisan et al., 2022) investigated interactive model cards, which extend traditional static cards with features for richer documentation exploration and direct model interaction. Drawing on insights from interviews with DL/ML experts and non-experts, their analysis highlights the importance of thoughtful design and interactivity, offering guidelines for more human-centered model cards.
@@
Meanwhile, some other investigators have focused on the development and distribution of the PTMs themselves. Han et al. (Han et al., 2021) comprehensively reviewed the past and present developments in PTMs. They analyzed recent advances driven by the availability of large training datasets and rapidly increasing computational power, and proposed several promising directions for future research. Jiang et al. (Jiang et al., 2022) first conducted a comprehensive empirical study in the pretrained model supply chain and analyzed eight popular model hubs that host PTMs and facilitate sharing between model developers and users. Their study revealed that the PTM supply chain plays a vital role in the DL supply chain and the existing security approaches for the PTM supply chain are insufficient.
@@
Furthermore, Zhang et al. (Zhang et al., 2023) introduced Model Spider to help users select appropriate pretrained models by computing and ranking similarity between token representations of potential models and the user’s intended features. Li et al. (Li et al., 2023) also proposed a similar approach for recommending appropriate PTMs to users, which we refer to as “Guided Recommendation” for brevity. They addressed the model selection challenge by formulating it as a model recommendation problem. Specifically, they leveraged related information, such as the features and labels of the models’ training datasets, along with their objective tasks, to compute guidance scores that help recommend suitable PTMs to users. Both approaches focus on selecting appropriate pretrained models for downstream tasks. However, Guided Recommendation emphasizes incorporating task-specific guidance to improve PTM recommendation, whereas Model Spider focuses on learning matching patterns between PTMs and tasks to enable effective selection. Nevertheless, the effectiveness of both Model Spider and Guided Recommendation depends on the availability of the pretrained models’ training datasets, the datasets for the downstream tasks and the downstream task objectives. In downstream application development scenarios, such datasets and precise task objectives are not always accessible to downstream application developers. As a result, these approaches are particularly well-suited for researchers and developers who have expertise in AI and are interested in exploring subtle differences among pretrained models, selecting models with the best performance across standardized evaluation criteria, and further fine-tuning them for specific application contexts.
@@
Moreover, Liu et al. (Liu et al., 2023) proposed MLTaskKG, a system designed to recommend machine learning (ML) and deep learning (DL) libraries or models to application developers based on the AI task specifications they provide. MLTaskKG constructs a comprehensive knowledge graph by aggregating publicly available pretrained models and libraries from diverse sources such as PapersWithCode (Meta AI, nd) and published literature, and modeling relationships between AI tasks, ML/DL models, and their underlying implementations. This knowledge graph is subsequently leveraged to recommend appropriate ML/DL libraries by matching developer-specified tasks, model characteristics, and implementation requirements. The approach is particularly advantageous for downstream application developers who have limited ML/DL expertise. However, MLTaskKG mainly emphasizes functional-oriented characteristics while paying limited attention to potential limitations of the recommended models or libraries, such as issues related to bias, safety, or license compatibility. Moreover, the project is no longer actively maintained and we are unable to replicate the system due to the lack of up-to-date implementation.
@@
3. Preparation of Candidate PTMs
The existing literature (Jiang et al., 2022, 2023, 2024) has examined several popular model hubs, such as Hugging Face (Hugging Face Company, nd), Model Zoo (Jing Yu Koh, nd), and ONNX Model Zoo (ONNX Model Zoo, nd), which serve as hosting platforms for publicly available pretrained models (PTMs). In this study, Hugging Face is selected as the model hosting platform, as it offers the largest, most diverse, and best-documented collection of pretrained models compared to other hubs. Notably, our approach can be easily extended to support PTMs on other sharing platforms with minor modifications, enabling PTMPicker to integrate models from alternative sources.
The component of “Model Collection” in Figure 3 demonstrates an example of the publication of a pretrained model on Hugging Face, highlighting key elements such as the model registry name, relevant tags, and the model card. When model developers publish a pretrained model on Hugging Face with a registry name, they typically specify a set of descriptive tags, such as task (i.e., the model’s designed purpose) and framework (i.e., the implementation platform). More importantly, the model publisher is also expected to provide a model card (a detailed specification of the model), which appears below the tags section. To streamline the creation of model cards, Hugging Face provides a standardized model card template (Hugging Face, nda) that can document key aspects, such as model function, training data, evaluation metrics, and interaction examples.
Figure 3. The workflow of  PTMPicker.
[…]
The workflow of experimental dataset preparation and PTMPicker.
The number of models hosted on Hugging Face is more than 544,000 as of the start date of our research. To efficiently scrape models hosted on Hugging Face, we resort to the popular Python web scraping framework, Scrapy (Zyte, etc., nd), which is an open-source and fast high-level web scraping and crawling framework licensed under BSD (Wikipedia, nda). Using the framework, we retrieved all available model registry entries and scraped all meaningful attributes for each model, including monthly downloads, user likes, and all other associated tags. In addition, we downloaded the associated model card by utilizing the huggingface_hub (Hugging Face community, nd) library for each of the collected model entries. In essence, the model card is a Markdown file (README.md) maintained in Git, and typically includes YAML front matter to describe metadata such as language, license, and task. Given the name of the model registry and the file extension (e.g..md), the library reliably downloads all .md files associated with the model, including the target model card.
@@ 3. Preparation of Candidate PTMs
4. PTMPicker
@@
4.1. Metadata Attributes Representation for Candidate Models
Inspired by the existing literature (Jiang et al., 2024; Schelter et al., 2017; Tsay et al., 2020; Li et al., 2022; Tsay et al., 2022), we compiled the metadata attributes proposed for PTM descriptions and adapted extraction scripts provided in these studies to process the downloaded model cards. Table 1 presents the complete set of metadata fields examined in this study. The metadata attributes drawn from the literature refer to those proposed in related studies, and we provide brief explanations for them [^3] Please refer to our artifact for a detailed explanation for each metadata. https://zenodo.org/records/15958951. In contrast, the metadata fields stemming from publicity tags are directly scraped from the model hosting platform and are generally considered self-explanatory. It is also worth noting that certain metadata fields may originate from both the Literature and the Publicity Tags. This is due to potential differences between attributes extracted from model cards and fields scraped directly from the model hosting publicity tags. With regard to duplicated metadata fields, we consolidated them into a single representation to avoid redundancy. In total, we have 33 distinct metadata attributes that constitute our predefined template for describing both models and developer-specified requirements.
[…]
Prompt design in the process of dataset preparation and model selection.
To effectively and efficiently extract metadata attributes from scraped model cards, we referred to an existing artifact (Jiang et al., 2024; Jiang et al., 2024), which achieves an attribute extraction accuracy of up to 94.39%. We adapted this artifact for the specific objectives of our research, as our work requires extracting a broader set of metadata attributes than the original implementation. Furthermore, we leveraged recent advances in large language models by incorporating a more reliable and capable LLM (i.e., GPT-4 (OpenAI, nd) in this study). Consequently, we were able to effectively extract the metadata attributes, as shown in the Metadata Extraction component of Figure 4. Listing 1 illustrates extracted metadata attributes for a pretrained model, including function, input_format, output_format, and task. In practice, the extraction process is implemented using the Python library Instructor (Developers, nd), a widely adopted tool for structuring the outputs of underlying LLMs, such as Mixtral (Mistral AI, nd) and GPT-4 (OpenAI, nd).
@@ 4.2. Model Search Requests Synthesis
To construct this dataset, we populated the predefined template with attribute values synthesized using large language models. The generation of these values follows mutation strategies based on prior work (Patil et al., 2023; Zhuo et al., 2024; Jia and Harman, 2010; Papadakis et al., 2019) and is guided by the metadata values extracted from the curated pretrained models. By leveraging the extracted attribute values and few-shot prompting, we designed effective prompts to generate diverse and semantically meaningful attribute values, as shown in the Search Request Synthesis component of Figure 4. For example, we can relax or tighten the license compatibility requirements, modify the input format (e.g., changing the input language or modality), or adjust the core functionality (e.g., shifting from text-to-image generation to image-to-text generation). Specifically, in terms of attributes with a limited value set, such as license, input_format, and output_format, we substitute their original values with semantically valid alternatives. Regarding attributes, such as function and limitation, we employ prompt-based LLMs with few-shot examples and mask-and-infill techniques to modify key semantic elements while preserving the fluency and coherence of the original descriptions. Listing 2 presents an example of model search request mutation. Harnessing the generative capabilities of large language models, we can effectively mutate given model-populated instances, thereby producing new, diverse, and semantically valid model search requests. Therefore, this approach not only broadens the scope of template-based model search requests but also improves the flexibility and adaptability of the search process, enabling it to accommodate more sophisticated model search demands.
@@ 4.3. Model Selection
Special attributes are those that require more nuanced handling and cannot be directly compared with the developer-specified attribute values, while trivial ones are all the remaining attributes, which do not need special consideration and can be used for comparison straightforwardly. For example, the trivial attribute function, which briefly summarizes the model’s primary purpose, can be directly used to perform an accurate text match between the function attribute in the model search request and that in the candidate model-populated instances. Specifically, for trivial attributes, the embedding similarity based on BM25 (Robertson et al., 2009), though not limited to it, is computed between the attribute values extracted from the candidate models and those specified in the model search requests. This semantic similarity comparison helps to achieve a more efficient matching process.
@@ 5.3. Evaluation Setup
Addressing these research questions requires both extracting metadata from model card descriptions and subsequently synthesizing structured model search requests using the extracted information. Thanks to rapid advances in large language models (Mistral AI, nd; OpenAI, nd), both metadata extraction and search request synthesis can be performed effectively and efficiently through well-designed prompts. In this study, we selected GPT-4 (OpenAI, nd) as our experimental LLM because of its state-of-the-art capabilities and widespread adoption in the software engineering research community. Specifically, we carefully designed the prompts shown in Figure 4 and provided three input-output examples to elicit the model to generate more concise and accurate results. We used the default 4k context window and retained all default GPT-4 settings except for the temperature and TopP parameters, which were manually set to 0.0 and 0.95, respectively. This configuration minimizes randomness in the output while preserving sufficient creativity to support effective extraction and generation of analysis results.

### v3

@@ 1. Introduction
PreTrained Models (PTMs) refer to neural network models that have been previously trained on large-scale datasets and saved for future reuse. In recent years, the existing literature on PTMs discusses the core issues and risks of the PTM supply chain and hubs (mitchell2019model; jiang2022empirical; castano2023exploring; kathikar2023assessing) and the detailed utilization of specific PTMs for special tasks (schmidhuber2015deep; antoniou2017data; dube2019automatic; xia2022systematic; qiu2020pre), such as molecular property prediction in chemistry. In addition to the literature on PTM supply chain risks and domain-specific applications, there is also a stream of research (bolya2021scalable; you2021logme; zhang2023model; liu2023task; davis2023reusing) dedicated to the identification and selection of suitable models. Specifically, Zhang et al. (zhang2023model) proposed Model Spider, which selects suitable pretrained models for specific tasks by computing and ranking the similarity between the candidate model token representation and the downstream task token representation. Model Spider better serves researchers and developers who are interested in the nuanced differences among various PTMs and seek to explore their performance across specific evaluation metrics. Liu et al. (liu2023task) introduces MLTaskKG to recommend ML/DL libraries/models for application developers based on the specifications of AI tasks specified by them. MLTaskKG primarily focuses on AI task descriptions while overlooking the limitations of recommended models or libraries, such as potential biases, safety concerns, and license compatibility.
The surge in PTMs has led to the emergence of Hugging Face (hfsite:2025; hfwiki:2025) as the de facto standard model hosting platform. Hugging Face, an online Git-based platform, hosted more than 543,000 diverse pretrained models at the time of our study, and the number of models continues to grow rapidly. This centralized repository could obviate the need for researchers to manually retrieve models from other heterogeneous hosting platforms. Although the platform provides model specifications, developers [^1] For simplicity, unless otherwise specified, the term developers hereafter refers to downstream application developers in this study. still face challenges in finding models that align with their requirements. In practice, application developers are often concerned not only with the intended functionality of potential target PTMs but also with aspects such as license compliance, model size, and interface usability. When searching using functional keywords or high-level task categories, an overwhelming number of candidate models are typically returned. Developers then need to devote considerable attention to manually and carefully examining and iterating all the candidate models, which is both labor-intensive and time-consuming.
@@ 2. Background & Motivation
Compared with training models from scratch, pretrained models (PTMs) substantially reduce cost and time requirements, and, when coupled with post-training compression, such as quantization—enable deployment on resource-constrained devices (egashira2024exploiting). This efficiency lowers barriers for developers by allowing goal-oriented reuse through inference with minimal integration effort, making PTM-centric workflows attractive across domains from general AI systems to embedded/IoT software (MSurvey9; MSurvey5). At the same time, widespread PTM adoption intersects with security considerations in which LLM-based components introduce new attack surfaces and must be assessed against evasion and supply-chain risks documented in recent surveys and empirical studies (MSurvey6; MSurvey7; MSurvey8; MTech1).
@@
The emergence and widespread adoption of PTMs has attracted significant research attention, with some researchers focusing on the development and analysis of model documentation, commonly referred to as model cards. Mitchell et al. (mitchell2019model) first proposed model cards for two types of Deep Learning (DL) models: one trained to identify smiling faces in images and the other trained to detect offensive comments in natural language texts. These model cards were intended to promote model transparency among not only users but also developers and stakeholders. Going one step further, Crisan et al. (crisan2022interactive) investigated interactive model cards, which extend traditional static cards with features for richer documentation exploration and direct model interaction. Drawing on insights from interviews with DL/ML experts and non-experts, their analysis highlights the importance of thoughtful design and interactivity, offering guidelines for more human-centered model cards.
@@
Meanwhile, some other investigators have focused on the development and distribution of the PTMs themselves. Han et al. (han2021pre) comprehensively reviewed the past and present developments in PTMs. They analyzed recent advances driven by the availability of large training datasets and rapidly increasing computational power, and proposed several promising directions for future research. Jiang et al. (jiang2022empirical) first conducted a comprehensive empirical study in the pretrained model supply chain and analyzed eight popular model hubs that host PTMs and facilitate sharing between model developers and users. Their study revealed that the PTM supply chain plays a vital role in the DL supply chain and the existing security approaches for the PTM supply chain are insufficient.
@@
Furthermore, Zhang et al. (zhang2023model) introduced Model Spider to help users select appropriate pretrained models by computing and ranking similarity between token representations of potential models and the user’s intended features. Li et al. (li2023guided) also proposed a similar approach for recommending appropriate PTMs to users, which we refer to as “Guided Recommendation” for brevity. They addressed the model selection challenge by formulating it as a model recommendation problem. Specifically, they leveraged related information, such as the features and labels of the models’ training datasets, along with their objective tasks, to compute guidance scores that help recommend suitable PTMs to users. Both approaches focus on selecting appropriate pretrained models for downstream tasks. However, Guided Recommendation emphasizes incorporating task-specific guidance to improve PTM recommendation, whereas Model Spider focuses on learning matching patterns between PTMs and tasks to enable effective selection. Nevertheless, the effectiveness of both Model Spider and Guided Recommendation depends on the availability of the pretrained models’ training datasets, the datasets for the downstream tasks and the downstream task objectives. In downstream application development scenarios, such datasets and precise task objectives are not always accessible to downstream application developers. As a result, these approaches are particularly well-suited for researchers and developers who have expertise in AI and are interested in exploring subtle differences among pretrained models, selecting models with the best performance across standardized evaluation criteria, and further fine-tuning them for specific application contexts.
@@
Moreover, Liu et al. (liu2023task) proposed MLTaskKG, a system designed to recommend machine learning (ML) and deep learning (DL) libraries or models to application developers based on the AI task specifications they provide. MLTaskKG constructs a comprehensive knowledge graph by aggregating publicly available pretrained models and libraries from diverse sources such as PapersWithCode (paperswithcode:2025) and published literature, and modeling relationships between AI tasks, ML/DL models, and their underlying implementations. This knowledge graph is subsequently leveraged to recommend appropriate ML/DL libraries by matching developer-specified tasks, model characteristics, and implementation requirements. The approach is particularly advantageous for downstream application developers who have limited ML/DL expertise. However, MLTaskKG mainly emphasizes functional-oriented characteristics while paying limited attention to potential limitations of the recommended models or libraries, such as issues related to bias, safety, or license compatibility. Moreover, the project is no longer actively maintained and we are unable to replicate the system due to the lack of up-to-date implementation.
@@
3. Preparation of Candidate PTMs
Figure 1. The workflow of PTMPicker.
The existing literature (jiang2022empirical; jiang2023ptmtorrent; jiang2024peatmoss) has examined several popular model hubs, such as Hugging Face (hfsite:2025), Model Zoo (modelzoo:2025), and ONNX Model Zoo (onnxmodel:2025), which serve as hosting platforms for publicly available pretrained models (PTMs). In this study, Hugging Face is selected as the model hosting platform, as it offers the largest, most diverse, and best-documented collection of pretrained models compared to other hubs. Notably, our approach can be easily extended to support PTMs on other sharing platforms with minor modifications, enabling PTMPicker to integrate models from alternative sources.
The component of “Model Collection” in Figure 1 demonstrates an example of the publication of a pretrained model on Hugging Face, highlighting key elements such as the model registry name, relevant tags, and the model card. When model developers publish a pretrained model on Hugging Face with a registry name, they typically specify a set of descriptive tags, such as task (i.e., the model’s designed purpose) and framework (i.e., the implementation platform). More importantly, the model publisher is also expected to provide a model card (a detailed specification of the model), which appears below the tags section. To streamline the creation of model cards, Hugging Face provides a standardized model card template (cardtemplate:2025) that can document key aspects, such as model function, training data, evaluation metrics, and interaction examples.
@@ 3. Preparation of Candidate PTMs
The number of models hosted on Hugging Face is more than 544,000 as of the start date of our research. To efficiently scrape models hosted on Hugging Face, we resort to the popular Python web scraping framework, Scrapy (scrapy:2025), which is an open-source and fast high-level web scraping and crawling framework licensed under BSD (bsd:2025). Using the framework, we retrieved all available model registry entries and scraped all meaningful attributes for each model, including monthly downloads, user likes, and all other associated tags. In addition, we downloaded the associated model card by utilizing the huggingface_hub (hfdownload:2025) library for each of the collected model entries. In essence, the model card is a Markdown file (README.md) maintained in Git, and typically includes YAML front matter to describe metadata such as language, license, and task. Given the name of the model registry and the file extension (e.g..md), the library reliably downloads all .md files associated with the model, including the target model card.
@@ 4. PTMPicker
4.1. Metadata Attributes Representation for Candidate Models
Inspired by the existing literature (jiang2024peatmoss; schelter2017automatically; tsay2020aimmx; li2022metadata; tsay2022extracting), we compiled the metadata attributes proposed for PTM descriptions and adapted extraction scripts provided in these studies to process the downloaded model cards. Table 1 presents the complete set of metadata fields examined in this study. The metadata attributes drawn from the literature refer to those proposed in related studies, and we provide brief explanations for them [^3] Please refer to our artifact for a detailed explanation for each metadata. https://zenodo.org/records/15958951. In contrast, the metadata fields stemming from publicity tags are directly scraped from the model hosting platform and are generally considered self-explanatory. It is also worth noting that certain metadata fields may originate from both the Literature and the Publicity Tags. This is due to potential differences between attributes extracted from model cards and fields scraped directly from the model hosting publicity tags. With regard to duplicated metadata fields, we consolidated them into a single representation to avoid redundancy. In total, we have 33 distinct metadata attributes that constitute our predefined template for describing both models and developer-specified requirements.
@@ 4.1. Metadata Attributes Representation for Candidate Models
To effectively and efficiently extract metadata attributes from scraped model cards, we referred to an existing artifact (mdextract:2025; jiang2024peatmoss), which achieves an attribute extraction accuracy of up to 94.39%. We adapted this artifact for the specific objectives of our research, as our work requires extracting a broader set of metadata attributes than the original implementation. Furthermore, we leveraged recent advances in large language models by incorporating a more reliable and capable LLM (i.e., GPT-4 (gpt4:2025) in this study). Consequently, we were able to effectively extract the metadata attributes, as shown in the Metadata Extraction component of Figure 4. Listing 1 illustrates extracted metadata attributes for a pretrained model, including function, input_format, output_format, and task. In practice, the extraction process is implemented using the Python library Instructor (pyinstructor:2025), a widely adopted tool for structuring the outputs of underlying LLMs, such as Mixtral (mixtral:2025) and GPT-4 (gpt4:2025).
@@ 4.2. Model Search Requests Synthesis
To construct this dataset, we populated the predefined template with attribute values synthesized using large language models. The generation of these values follows mutation strategies based on prior work (patil2023gorilla; zhuo2024bigcodebench; jia2010analysis; papadakis2019mutation) and is guided by the metadata values extracted from the curated pretrained models. By leveraging the extracted attribute values and few-shot prompting, we designed effective prompts to generate diverse and semantically meaningful attribute values, as shown in the Search Request Synthesis component of Figure 4. For example, we can relax or tighten the license compatibility requirements, modify the input format (e.g., changing the input language or modality), or adjust the core functionality (e.g., shifting from text-to-image generation to image-to-text generation). Specifically, in terms of attributes with a limited value set, such as license, input_format, and output_format, we substitute their original values with semantically valid alternatives. Regarding attributes, such as function and limitation, we employ prompt-based LLMs with few-shot examples and mask-and-infill techniques to modify key semantic elements while preserving the fluency and coherence of the original descriptions. Listing 2 presents an example of model search request mutation. Harnessing the generative capabilities of large language models, we can effectively mutate given model-populated instances, thereby producing new, diverse, and semantically valid model search requests. Therefore, this approach not only broadens the scope of template-based model search requests but also improves the flexibility and adaptability of the search process, enabling it to accommodate more sophisticated model search demands.
@@ 4.3. Model Selection
Special attributes are those that require more nuanced handling and cannot be directly compared with the developer-specified attribute values, while trivial ones are all the remaining attributes, which do not need special consideration and can be used for comparison straightforwardly. For example, the trivial attribute function, which briefly summarizes the model’s primary purpose, can be directly used to perform an accurate text match between the function attribute in the model search request and that in the candidate model-populated instances. Specifically, for trivial attributes, the embedding similarity based on BM25 (robertson2009probabilistic), though not limited to it, is computed between the attribute values extracted from the candidate models and those specified in the model search requests. This semantic similarity comparison helps to achieve a more efficient matching process.
@@ 5.3. Evaluation Setup
Addressing these research questions requires both extracting metadata from model card descriptions and subsequently synthesizing structured model search requests using the extracted information. Thanks to rapid advances in large language models (mixtral:2025; gpt4:2025), both metadata extraction and search request synthesis can be performed effectively and efficiently through well-designed prompts. In this study, we selected GPT-4 (gpt4:2025) as our experimental LLM because of its state-of-the-art capabilities and widespread adoption in the software engineering research community. Specifically, we carefully designed the prompts shown in Figure 4 and provided three input-output examples to elicit the model to generate more concise and accurate results. We used the default 4k context window and retained all default GPT-4 settings except for the temperature and TopP parameters, which were manually set to 0.0 and 0.95, respectively. This configuration minimizes randomness in the output while preserving sufficient creativity to support effective extraction and generation of analysis results.